OpenStudy (anonymous):

@cody7744 MEDAL AND FAN How many solutions does the equation have? 4x + 1 = 5 + 2(2x - 4) A. One Solution B. Infinite Solutions C. No Solution

OpenStudy (anonymous):

4x+1=5+2(2x-4) 4x+1=5+4x-8

OpenStudy (anonymous):

c

OpenStudy (anonymous):

this one is hard

OpenStudy (anonymous):

Which expressions would complete this equation so that it has one solution? 4(2x – 3) – 4x = Choose exactly two answers that are correct. A. 4x – 12 B. 3x – 4 C. 3(2x – 6) D. 4(x – 5)

OpenStudy (anonymous):

A - B - C - D -

OpenStudy (anonymous):

@cody7744

OpenStudy (anonymous):

c

OpenStudy (anonymous):

and b

OpenStudy (anonymous):

thx